OpenStudy (anonymous):

A number cube has faces labeled 1 through 6. What is the probability of the event that is complementary to the event of rolling a 6 on the number cube?

OpenStudy (kc_kennylau):

Probability of rolling a 6 = 1/6

OpenStudy (kc_kennylau):

Complementary = 1 - 1/6 = 5/6

OpenStudy (anonymous):

what is the answer

OpenStudy (kc_kennylau):

5/6

OpenStudy (anonymous):

i have more mquestion can u help...

OpenStudy (anonymous):

i will post it